OpenStudy (anonymous):

PLEASE HELPP ASAP 1. suppose you invest $1,500 at an annual interest rate of 6.9% compounded continuously. how much will you have in the account after 10 years A. $14,952.87 B. $29,905.73 C. $2,990.57 D. $1,607.15 2. what is the value of log4 (2) A. 2/3 B. 1/4 C. 1/3 D. 1/2 3. solve log(2x+1)=3 A. 999 B. 500 C. 999/2 D. 1/2 4. solve the equation log(x+5)-logx=3 A. 0.0505 B. 0.0050 C. 2.5 D. 199.8

OpenStudy (anonymous):

2: 1/2 3: \[\frac{ \log(x+1) }{ \log(2) }\] 4: I got 0.26198

OpenStudy (anonymous):

i dont understand how you got those answers for 2 and 4 because the answers isnt on the choices

OpenStudy (anonymous):

yeah, That's what I was thinking.

OpenStudy (mathmale):

Regarding #4: Use rules of logs (and of exponentiation) to simplify log(x+5)-log(x) = 3. Hint: How do we write log a - log b as a single logarithm?

OpenStudy (anonymous):

see thats what im confused about :P

OpenStudy (mathmale):

Here are the relevant rules of logs: (a) log a + log b = log a*b (product of a and b) (b) log a - log b = log a/b (quotient of a and b) (c) log a^b = b*log a Really important to know these 3 rules and how to apply them. For practice: re-write log 2 - log 3 as a single log.

OpenStudy (mathmale):

2. what is the value of log4 (2)? RockStarGirl: Before we start on this Problem #2, let's see whether we can agree on how the problem statement reads. When I see "what is the value of log4 (2)," I can't help but wonder whether this means "the logarithm to the base 4 of 2."

OpenStudy (mathmale):